OpenStudy (anonymous):

simplify ln9-ln4ln3. i have ln9-ln81 so far

hartnn (hartnn):

use \(\huge ln A-lnB=ln(A/B)\)

Parth (parthkohli):

\[\ln 9 - \ln 81 = \ln{9 \over 81}\]

Parth (parthkohli):

@hartnn won. :P

hartnn (hartnn):

i wasn't in a race

OpenStudy (anonymous):

nice one cheers guys!!

hartnn (hartnn):

that can still be simplified, know how ?

OpenStudy (anonymous):

yeh think so

Parth (parthkohli):

Cool.

hartnn (hartnn):

so tell us what u get finally, so that we can verify it for u.

OpenStudy (anonymous):

i get 0.1 is that right?

hartnn (hartnn):

9/81 = ?

hartnn (hartnn):

and no, thats incorrect

OpenStudy (anonymous):

yeh i got 9/81 which is 1/9 then 0.1

hartnn (hartnn):

ok, don't convert it in decimal, secondly, u forgot to write ln now u know how to simplify ln(1/9) ?

hartnn (hartnn):

1/9 = \(9^{-1}\)

OpenStudy (anonymous):

ok thank you

hartnn (hartnn):

so finally u get -ln9 = -ln 3^2 = -2ln3

OpenStudy (anonymous):

yeh got it now thanks for help much appreciated

hartnn (hartnn):

welcome ^_^